OpenStudy (anonymous):

1 over x squared plus 8 over x plus 15 = 0

OpenStudy (anonymous):

x= -1/3 , x= -1/5

OpenStudy (gorv):

\[\frac{ 1 }{ x^2}+\frac{ 8 }{ x+15}=0\]

OpenStudy (gorv):

is it your Q??

OpenStudy (anonymous):

I think what you just posted was the question and he wants the solutions @gorv

OpenStudy (gorv):

yeah that what i asked is it your Q ..just read above your last comment

OpenStudy (anonymous):

How

OpenStudy (anonymous):

How what?

OpenStudy (anonymous):

How did you do it

OpenStudy (anonymous):

wolframalpha , I just typed in the equation

OpenStudy (anonymous):

I don't really know how to do the problem, you'll have to get someone else to show you how to do it, I'm sorry

OpenStudy (anonymous):

Okay